OpenStudy (anonymous):

An interior angle and an exterior angle are adjacent. The exterior angle measures 113°. What is the measure of the interior angle? 67° 77° 63° 73°

OpenStudy (anonymous):

remember that an interior angle plus an exterior angle = 180 degrees

OpenStudy (anonymous):

so x+113 = 180 x = 180-113 x = 63

OpenStudy (anonymous):

*67

OpenStudy (anonymous):

not 63, sorry

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

so its 67??